# Vector Quantized-Elites: Unsupervised and Problem-Agnostic Quality-Diversity Optimization

### v2

@@ V Experiments
For our experiments, we selected two state-covering tasks using a mobile robot, two pose-reaching tasks with a 7 degrees of freedom robotic arm, and a hard exploration and discrete task inspired by the MiniGrid Reinforcement Learning environment [36]. These tasks were chosen to showcase the versatility and effectiveness of VQ-Elites in solving problems without requiring prior knowledge of the environment or the system. Moreover, these experiments demonstrate the broader potential of VQ-Elites beyond these specific scenarios.
@@ V-A Mobile Robot Experiments
For the space-covering experiments, we employ a differential drive mobile robot. The robot is deployed in two environments, one without obstacles and one with a L-shaped environment as depicted in Fig. 2. During the evolution phase, the algorithm evolves a population that contains the weights of a NN, which acts as the policy. This policy takes as input the state of the robot $\mathbf{s}=\{x,y,\cos(\theta),\sin(\theta)\}$, and produces a $\mathbf{v}=\{v_{L},v_{R}\}$ that includes the velocity that should be applied on each wheel. The goal of this task is for the robot to cover the whole environment while reaching each final position as fast as possible. Hence the fitness function is:
$f=-\sum_{i}^{N}\|\boldsymbol{p}_{i}-\boldsymbol{p}_{N}\|_{2},$ (8)
@@ V-C MiniGrid Exploration
The MiniGrid experiment (Fig. 3) consists of a grid map, and an agent moving (with discrete actions) inside it. We use this experiment to investigate the performance of our approach in a domain other than robotics. The ultimate goal of the task is to discover individuals through the evolution process to traverse all possible tiles of the map. The agent has a field of view (FOV) of 3 tiles and can perform 5 actions, $\mathcal{A}=${up, right, down, left, stay still}, which are selected by the policy. For the policy’s input, we use snapshots of the current agent’s position while respecting its FOV, along with the list of keys that have been gathered using one-hot encoding. Hence, these snapshots contain information in pixels inside the FOV of the agent, while the rest are masked out. Additionally, the snapshot of the last step in the environment also serves as our raw BD. The size of the snapshots is $18\times 18$.
@@ V-D Evaluation Metrics
To evaluate the performance of our algorithm, we employ most of the metrics used in AURORA [9], but we also introduce two new metrics. Specifically, we evaluate our approach in terms of coverage, grid mean fitness, QD score, and our newly proposed Effective Diversity Ratio and Coverage Diversity Score.
@@ V-D1 Coverage
V-D2 Quality-Diversity Score
The QD score is used to assess the performance of an algorithm in terms of both the diversity and the quality of solutions in the BD space. It is calculated as the sum of the fitness values of all the individuals across the grid. By combining diversity (coverage of the grid) and quality (fitness of solutions), the QD score provides a single, unified measure to evaluate the effectiveness of a given algorithm. Higher QD scores indicate better exploration and exploitation of the BD space.
V-D3 Grid Mean Fitness
This metric assesses the overall quality of a container by calculating the mean fitness of individuals in a grid of the hand-coded BD space. For each cell, only the individual with the highest fitness is retained, and their fitnesses are averaged to compute the grid mean fitness. This approach ensures that the metric is independent of the solution distribution within the BD space, enabling fair comparisons across algorithms using learned BD spaces. This metric is a normalized version of the QD score, and is basically a qualitative visual metric. As such we apply it only in the mobile robot and MiniGrid scenarios.
V-D4 Effective Diversity Ratio
@@
This metric reflects how effectively the algorithm explores the behavior outcome space, discounting redundancy among individuals whose executions lead to similar behaviors. A higher EDR indicates broader behavior coverage and less repetition in the discovered solutions.
V-D5 Coverage Diversity Score
We also introduce the Coverage Diversity Score (CDS), a metric designed to assess behavioral diversity in the context of archive coverage. It is defined as the product of Coverage and EDR:
@@ V-D5 Coverage Diversity Score
While EDR captures the raw diversity of the archive, it can be misleading in edge cases. For example, at the very beginning of evolution, an archive containing a single individual will yield an EDR of $100\%$, despite offering no meaningful diversity. Conversely, as the archive grows, the EDR may become artificially low even if the behavior spread remains satisfactory. By incorporating coverage, the CDS mitigates these issues, providing a more balanced view of diversity relative to the size of the explored behavior space. It offers a more robust indicator of how effectively an algorithm populates the behavior space with both broad and distinct solutions.
@@ VI Results
In this section, we showcase the results of our algorithm in the experimental evaluation, starting with the mobile robot experiments, then continuing to robotic arm experiments and the MiniGrid exploration task, and finally presenting ablation studies we performed on our algorithm.
@@ VI-A Mobile Robot Results
For the mobile robot experiments, we compare VQ-Elites algorithm with the classic MAP-Elites algorithm, the AURORA approach, as well as the AURORA after integrating our two proposed components, behavior space bounding and cooperation, which from now on we will refer to as AURORA. We should mention that a variation of this experiment is also present in the AURORA paper [9], hence it is ideal for comparison purposes, however, we are unable to exactly replicate the authors’ results [^2] We used the exact hyperparameters presented in the original paper and communicated with the authors., as demonstrated in the manuscript.
In the case where the robot operates in an obstacle-free environment, we can observe that all the algorithms, except the vanilla AURORA, perform almost perfectly in terms of coverage, assigning solutions to nearly every grid cell (Fig. 4). In terms of the QD score, we see that the classic MAP-Elites algorithm produces more high-quality and diverse solutions compared to the other approaches. Both of these observations are expected, since we can easily define a behavior grid manually, and still every cell will represent a feasible behavior. Nonetheless, it is evident that algorithms that assume no prior knowledge about the environment, such as ours and other unsupervised QD algorithms infused with our cooperation and behavior space bounding components, can achieve the performance of MAP-Elites without full knowledge of the task (answers to Question 1). Additionally, considering the EDR and CDS metrics, the VQ-Elites is producing slightly more diverse behaviors compared to MAP-Elites and AURORA. This indicates that despite the coverage of VQ-Elites being lower compared to the other two, the behaviors produced do a better job of promoting diversity while maintaining a high-quality archive.
@@
Transitioning to the L-shaped environment, our approach and AURORA outperform the others by a large margin in terms of coverage, while VQ-Elites exhibits the best QD score of all (Fig. 5). This is one type of task that makes the existence of unsupervised QD algorithms essential, since, if we do not possess full knowledge about the task at hand, and particularly in this scenario, we do not know the environment, we cannot construct a grid such that every cell corresponds to a possible behavior (answers to Question 2). On the other hand, a data-driven unsupervised approach is more than capable of constructing such a grid. Although MAP-Elites has low coverage, it outperforms the other algorithms in the effective diversity metric. This is expected since the hand-crafted behavior grid is well-discretized, enabling the algorithm to provide a small but diverse archive. However, this is not the case in the CDS metric, where VQ-Elites has achieved a higher score since the coverage is taken into account. Additionally, between VQ-Elites and AURORA, our algorithm performs better in the EDR and CDS metrics, since it does not rely on an unstructured archive, which populates the archive based on heuristics that may result in many duplicate behaviors.
@@
Lastly, since the experiment is two-dimensional, we also provide plots, which are hard or infeasible to do or comprehend in high dimensions. These plots depict the grid mean fitness, final archives of the algorithms, and the behavior descriptors for each algorithm per scenario in Fig. 6. In the first column of each plot, we illustrate the grid mean fitness by discretizing the robot’s operating space into an $8\times 8$ grid, and the color of each cell represents the normalized fitness of the best solution assigned to each cell. The second column showcases the actual solutions VQ-Elites produced, while the last one depicts the BDs of the solutions that emerged. Overall, we observe that VQ-Elites and AURORA are able to effectively explore the behavior space, produce meaningful latent representations, and in the case of VQ-Elites, it automatically performs clustering of the behaviors, constructing a meaningful behavior space grid.
Figure 6: Visualizations that demonstrate the mobile experiment’s outcomes of the three algorithms in the two different environments. Left: Obstacle-free experiment outcome. Right: L-Shaped experiment outcome. In both plots, the first column indicates the grid mean fitness in a $8\times 8$ grid, while the second depicts where every learned behavior ends up in the environment. The third column is a visualization of the behavior descriptors that describe these behaviors.
@@ VI-B Robotic Arm Results
For the first scenario, it is evident that MAP-Elites is the best of the batch in both coverage and QD score. Then VQ-Elites and AURORA follow, while VQ-Elites is slightly better. This fact indicates that the way an unstructured archive is updated is highly likely to discard high-quality and diverse individuals in order to satisfy the minimum distance between individuals rule (answers to Question 1). Out of the four algorithms, vanilla AURORA performs the worst.
The second case resembles the mobile robot in the L-shaped environment. However, in this task, the constraint is on the robot itself, by reducing some joint limits. Therefore, we change the limits of joints 1 and 2 (see Fig. 9) from [2.967,2.967][-2.967,2.967] to [0.5,0.5][-0.5,0.5]. Despite that, our algorithm has a completely agnostic nature due to the unsupervised learning model, and hence, it outperforms all the other approaches by a large margin in both coverage and QD score (Fig. 8). We should note that this is highly predictable because in order for the MAP-Elites to achieve a competitive performance, it requires knowledge of the constraints to construct a proper behavior descriptor grid. Lastly, both AURORA and vanilla AURORA perform similarly to the MAP-Elites algorithm despite its unsupervised nature, while the QD score of the vanilla variation is the worst of the three due to the way the archive is constructed, as noted previously (answers to Question 2). Focusing on the effective diversity and coverage diversity score, our algorithm is still outperforming the others. For instance, the $80\%$ of VQ-Elites’s archive is filled with unique behaviors in the former metric while achieving a near-perfect score in the latter. Surprisingly, the AURORA and vanilla AURORA fail to provide a competent amount of diverse behaviors.
@@
Figure 8: Results for the robotic arm pose reaching task with system constraints (i.e. limitation in some joints) for VQ-Elites, MAP-Elites with hard-coded BDs, and AURORA. Each graph line represents the median and the 25th and 75th quantiles over ten runs for the corresponding algorithm.
@@ VI-C MiniGrid Results
In the MiniGrid experiment, we evaluate the performance of the algorithms in a discrete and challenging exploration setting. The results, shown in Fig. 11, reveal that convergence in terms of coverage is significantly slower than in previous tasks. Notably, while VQ-Elites converges more slowly than the other methods, it achieves near-perfect effective diversity and coverage diversity (answers to Question 5 and 6). In contrast, both variants of the AURORA algorithm converge faster but fail to produce a truly diverse archive.
@@
Figure 10: Mean Grid Fitness of the MiniGrid exploration task for VQ-Elites, MAP-Elites with hard-coded BDs, and AURORA
@@ VI-D Ablation Study
Furthermore, we investigate the importance of the cooperation component on our algorithms as well as on the AURORA algorithm. As stated, VQ-Elites is quite robust that the cooperation does not have any impact on its performance at convergence, since the VQ-VAE model, eventually, converges to an optimal grid structure (answers to Question 4). As expected, while the cooperation mechanism is active, the QD score may fluctuate or remain low, since the algorithm intentionally allows lower-quality solutions to enter the archive in favor of promoting diversity. On the other hand, the AURORA algorithm performs poorly without this component (Fig. 13).
@@ VII Discussion and Future Work
Another limitation concerns the scalability of the approach to very high-dimensional behavior spaces, especially when the raw descriptors are unstructured. While VQ-Elites effectively compresses these spaces using a learned representation, the VQ-VAE model may struggle to disentangle meaningful features in complex environments without careful design. This limitation could potentially be mitigated by incorporating explicit structure into the latent space. For instance, enforcing Riemannian [37] or $SO(3)$ manifold constraints may lead to more meaningful and interpretable representations for behaviors with geometric or rotational properties. Such structure could be introduced either through architectural constraints or auxiliary objective functions during training.
@@
Moreover, while VQ-Elites was designed to be general and task-agnostic, this generality may come at the cost of suboptimality in highly structured domains. Future research could explore hybrid designs where partial task knowledge or hand-crafted descriptors are used to guide the model in early phases, later transitioning to fully learned representations as more data becomes available. Additionally, techniques such as attention mechanisms [38] or multi-modal inputs (e.g., fusing proprioceptive and visual data) could enhance representation quality in challenging domains.
@@
Finally, in the future, we want to incorporate generative modeling techniques such as diffusion [39, 40], or flow-matching [41], in order to model the distribution of solutions and generate more high-quality and diverse solutions. An ideal outcome would be to use VQ-Elites for finding the minimum amount of solutions, and then via the generative modeling techniques, we would able to fill the archive successfully without being limited by the run times of the evolutionary algorithm and its sequential nature. Lastly, another interesting direction would be an epistemic uncertainty-guided evolution frame. The intuition here lies in the fact that epistemic uncertainty decreases when the number of data we train our model increases [42], hence, we can use the epistemic uncertainty of the model as a way to guide the QD evolution into more uncertain behaviors, with the goal of reducing the epistemic uncertainty of the VQ-VAE model, and consequently creating the most optimal behavioral space possible. For this purpose, we can borrow ideas from surrogate-assisted QD algorithms [43].

### v3

@@ V Experiments
For our experiments, we selected two state-covering tasks using a mobile robot, two pose-reaching tasks with a 7 degrees of freedom robotic arm, and a hard exploration and discrete task inspired by the MiniGrid Reinforcement Learning environment [chevalier2018minimalistic]. These tasks were chosen to showcase the versatility and effectiveness of VQ-Elites in solving problems without requiring prior knowledge of the environment or the system. Moreover, these experiments demonstrate the broader potential of VQ-Elites beyond these specific scenarios.
@@ V-A Mobile Robot Experiments
For the space-covering experiments, we employ a differential drive mobile robot. The robot is deployed in two environments, one without obstacles and one with a L-shaped environment as depicted in Fig. 2. During the evolution phase, the algorithm evolves a population that contains the weights of a NN, which acts as the policy. This policy takes as input the state of the robot $\mathbf{s}=\{x,y,\cos(\theta),\sin(\theta)\}$, and produces a $\mathbf{v}=\{v_{L},v_{R}\}$ that includes the velocity that should be applied on each wheel. The goal of this task is for the robot to cover the whole environment while reaching each final position as fast as possible. Hence, the fitness function is:
$f=\sum_{i}^{N}e^{(-\|\boldsymbol{p}_{i}-\boldsymbol{p}_{N}\|_{2})},$ (8)
@@ V-C MiniGrid Exploration
The MiniGrid experiment (Fig. 3) consists of a grid map, and an agent moving (with discrete actions) inside it. We use this experiment to investigate the performance of our approach in a domain other than robotics. The ultimate goal of the task is to discover individuals through the evolution process to traverse all possible tiles of the map. The agent has a field of view (FOV) of 3 tiles and can perform 5 actions, $\mathbf{A}=${up, right, down, left, stay still}, which are selected by the policy. For the policy’s input, we use snapshots of the current agent’s position while respecting its FOV, along with the list of keys that have been gathered using one-hot encoding. Hence, these snapshots contain information in pixels inside the FOV of the agent, while the rest are masked out. Additionally, the snapshot of the last step in the environment also serves as our raw BD. The size of the snapshots is $18\times 18$.
@@ V-D Evaluation Metrics
To evaluate the performance of our algorithm, we employ most of the metrics used in AURORA [aurora], but we also introduce two new metrics. Specifically, we evaluate our approach in terms of coverage, (Projected) QD score, and our newly proposed Effective Diversity Ratio and Coverage Diversity Score. We also introduce a visualization method for the unsupervised setting, where the algorithms’ archives are projected onto a grid defined by the ground-truth behavior descriptor. We refer to this visualization as the Elite Grid Projection.
@@ V-D1 Coverage
V-D2 Projected Quality-Diversity Score
The QD score is used to assess the performance of an algorithm in terms of both the diversity and the quality of solutions in the BD space. It is calculated as the sum of the fitness values of all the individuals across the grid. By combining diversity (coverage of the grid) and quality (fitness of solutions), the QD score provides a single, unified measure to evaluate the effectiveness of a given algorithm. Higher QD scores indicate better exploration and exploitation of the BD space. However, for the QD score to be meaningful, it requires all the algorithms to have the same behavior space grid, which is impossible in an unsupervised setting. Hence, we use the Projected QD (PQD) score, which is a proxy of the QD score. The PQD score is calculated by discretizing the ground truth behavior space into cells, for each cell, we keep the elite, and a new projected archive is generated. Then, by calculating the classic QD score on this newly generated projected archive, we get the PQD score.
V-D3 Effective Diversity Ratio
@@
This metric reflects how effectively the algorithm explores the behavior outcome space, discounting redundancy among individuals whose executions lead to similar behaviors. A higher EDR indicates more unique behaviors and less redundancy in the discovered solutions.
V-D4 Coverage Diversity Score
We also introduce the Coverage Diversity Score (CDS), a metric designed to assess behavioral diversity in the context of archive coverage in the unsupervised setting. It is defined as the product of Coverage and EDR:
@@ V-D4 Coverage Diversity Score
While EDR represents the inverse redundancy in the archive, it does not always correspond directly to meaningful diversity. For example, at the very beginning of evolution, an archive containing a single individual will yield an EDR of $100\%$, despite offering no meaningful diversity. Conversely, as the archive grows, the EDR may become artificially low even if the behavior spread remains satisfactory. By incorporating coverage, the CDS score mitigates these issues, providing a more balanced view of diversity relative to the size of the explored behavior space. It offers a more robust indicator of how effectively an algorithm populates the behavior space with both broad and distinct solutions.
@@
Evaluation Hyperparameter
Projection Grid Cells $30\times 30$ 400 $18\times 18$
@@ VI Results
In this section, we showcase the results of our algorithm in the experimental evaluation, starting with the mobile robot experiments, then continuing to robotic arm experiments and the MiniGrid exploration task, and finally presenting ablation studies we performed on our algorithm. We report all the hyperparameters used in each setting at Table I. For the AURORA-based methods, we used the same environment-specific hyperparameters (target archive size, population, etc.) as for VQ-Elites, and followed the values of the original paper for the rest hyperparameters [aurora] [^2] We also refer to the supplementary material for the complete list of hyperparameters..
@@ VI-A Mobile Robot Results
For the mobile robot experiments, we compare VQ-Elites algorithm with the classic MAP-Elites algorithm, the AURORA approach, as well as the AURORA after integrating our two proposed components, behavior space bounding and cooperation, which from now on we will refer to as AURORA. We should mention that a variation of this experiment is also present in the AURORA paper [aurora], hence it is ideal for comparison purposes, however, we are unable to exactly replicate the authors’ results [^3] We used the exact hyperparameters presented in the original paper and communicated with the authors., as demonstrated in the manuscript.
In the case where the robot operates in an obstacle-free environment, we can observe that all the algorithms, except the vanilla AURORA, perform almost perfectly in terms of coverage, assigning solutions to nearly every grid cell (Fig. 4). In terms of the PQD score, we see that the classic MAP-Elites algorithm produces more high-quality and diverse solutions compared to the other approaches. Both of these observations are expected, since we can easily define a behavior grid manually, and still every cell will represent a feasible behavior. Nonetheless, it is evident that algorithms that assume no prior knowledge about the environment, such as ours and other unsupervised QD algorithms infused with our cooperation and behavior space bounding components, can achieve the performance of MAP-Elites without full knowledge of the task (answers to Question 1). Additionally, considering the EDR and CDS metrics, VQ-Elites is producing new behaviors more effectively compared to AURORA and it is on par with MAP-Elites. This indicates that despite the coverage of VQ-Elites being lower compared to the other two, the behaviors produced do a better job of promoting diversity while maintaining a high-quality archive.
@@
Transitioning to the L-shaped environment, our approach and AURORA outperform the others by a large margin in terms of coverage, while VQ-Elites exhibits the best PQD score of all (Fig. 5). This is one type of task that makes the existence of unsupervised QD algorithms essential, since, if we do not possess full knowledge about the task at hand, and particularly in this scenario, we do not know the environment, we cannot construct a grid such that every cell corresponds to a possible behavior (answers to Question 2). On the other hand, a data-driven unsupervised approach is more than capable of constructing such a grid. Although MAP-Elites has low coverage, it outperforms the other algorithms in the effective diversity metric. This is expected since the hand-crafted behavior grid is well-discretized, enabling the algorithm to provide a small but diverse archive. However, this is not the case in the CDS metric, where VQ-Elites has achieved a higher score since the coverage is taken into account. Additionally, between VQ-Elites and AURORA, our algorithm performs better in the EDR and CDS metrics, since it does not rely on an unstructured archive, which populates the archive based on heuristics that may result in many duplicate behaviors.
@@
Lastly, since the experiment is two-dimensional, we also provide plots, which are hard or infeasible to do or comprehend in high dimensions. These plots depict the elite grid projection, final archives of the algorithms, and the behavior descriptors for each algorithm per scenario in Fig. 6. In the first column of each plot, we illustrate the elite grid projection by discretizing the robot’s operating space into an $8\times 8$ grid, and the color of each cell represents the normalized fitness of the best solution assigned to each cell. The second column showcases the actual solutions VQ-Elites produced, while the last one depicts the BDs of the solutions that emerged. Overall, we observe that VQ-Elites and AURORA are able to effectively explore the behavior space, produce meaningful latent representations, and in the case of VQ-Elites, it automatically performs clustering of the behaviors, constructing a meaningful behavior space grid.
Figure 6: Visualizations that demonstrate the mobile experiment’s outcomes of the three algorithms in the two different environments. Left: Obstacle-free experiment outcome. Right: L-Shaped experiment outcome. In both plots, the first column indicates the elite grid projection, where we discretize the solution space in an $8\times 8$ grid and keep the fitness of the elite individual. The second depicts where every learned behavior ends up in the environment. The third column is a visualization of the behavior descriptors that describe these behaviors.
@@ VI-B Robotic Arm Results
For the first scenario, it is evident that MAP-Elites is the best of the batch in both coverage and PQD score. VQ-Elites and AURORA follow, while VQ-Elites is slightly better (Fig. 8). This fact indicates that the way an unstructured archive is updated is highly likely to discard high-quality and diverse individuals in order to satisfy the minimum distance between individuals rule (answers to Question 1). Out of the four algorithms, vanilla AURORA performs the worst.
@@
The second case resembles the mobile robot in the L-shaped environment. However, in this task, the constraint is on the robot itself, by reducing some joint limits. Therefore, we change the limits of joints 1 and 2 (see Fig. 7) from [2.967,2.967][-2.967,2.967] to [0.5,0.5][-0.5,0.5]. Despite that, our algorithm has a completely agnostic nature due to the unsupervised learning model, and hence, it outperforms all the other approaches by a large margin in both coverage and PQD score (Fig. 9). We should note that this is highly predictable because in order for the MAP-Elites to achieve a competitive performance, it requires knowledge of the constraints to construct a proper behavior descriptor grid. Lastly, both AURORA and vanilla AURORA perform similarly to the MAP-Elites algorithm despite their unsupervised nature, while the PQD score of the vanilla variation is the worst of the three due to the way the archive is constructed, as noted previously (answers to Question 2). Focusing on the effective diversity and coverage diversity score, our algorithm is still outperforming the others. For instance, the $50\%$ of VQ-Elites’s archive is filled with unique behaviors in the former metric while achieving the best score over all algorithms in the latter. Surprisingly, the AURORA and vanilla AURORA fail to provide a competent amount of diverse behaviors.
Figure 9: Results for the robotic arm pose reaching task with system constraints (i.e. limitation in some joints) for VQ-Elites, MAP-Elites with hard-coded BDs, AURORA, and AURORA. Each graph line represents the median and the 25th and 75th quantiles over ten runs for the corresponding algorithm.
@@ VI-C MiniGrid Results
In the MiniGrid experiment, we evaluate the performance of the algorithms in a discrete and challenging exploration setting. The results reveal that convergence of VQ-Elites in terms of coverage is significantly slower than in previous tasks (Fig. 1011), but it outperforms by quite some margin the other unsupervised algorithms on the effective diversity and coverage diversity metrics (answers to Questions 5 and 6). In contrast, both variants of the AURORA algorithm converge faster but fail to produce a truly diverse archive.
@@
Figure 11: Elite Grid Projection of the MiniGrid exploration task for VQ-Elites, MAP-Elites with hard-coded BDs, and AURORA
@@ VI-D Ablation Study
Furthermore, we investigate the importance of the cooperation component on our algorithms as well as on the AURORA algorithm. As stated, VQ-Elites is quite robust that the cooperation does not have any impact on its performance at convergence, since the VQ-VAE model, eventually, converges to an optimal grid structure (answers to Question 4). As expected, while the cooperation mechanism is active, the PQD score may fluctuate or remain low, since the algorithm intentionally allows lower-quality solutions to enter the archive in favor of promoting diversity. On the other hand, the AURORA algorithm performs poorly without this component (Fig. 13).
@@ VII Discussion and Future Work
Another limitation concerns the scalability of the approach to very high-dimensional behavior spaces, especially when the raw descriptors are unstructured. While VQ-Elites effectively compresses these spaces using a learned representation, the VQ-VAE model may struggle to disentangle meaningful features in complex environments without careful design. This limitation could potentially be mitigated by incorporating explicit structure into the latent space. For instance, enforcing Riemannian [lin2008riemannian] or $SO(3)$ manifold constraints may lead to more meaningful and interpretable representations for behaviors with geometric or rotational properties. Such structure could be introduced either through architectural constraints or auxiliary objective functions during training.
@@
Moreover, while VQ-Elites was designed to be general and task-agnostic, this generality may come at the cost of suboptimality in highly structured domains. Future research could explore hybrid designs where partial task knowledge or hand-crafted descriptors are used to guide the model in early phases, later transitioning to fully learned representations as more data becomes available. Additionally, techniques such as attention mechanisms [vaswani2017attention] or multi-modal inputs (e.g., fusing proprioceptive and visual data) could enhance representation quality in challenging domains.
@@
Finally, in the future, we want to incorporate generative modeling techniques such as diffusion [ho2020denoising, yang2023diffusion], or flow-matching [lipman2022flow], in order to model the distribution of solutions and generate more high-quality and diverse solutions. An ideal outcome would be to use VQ-Elites for finding the minimum amount of solutions, and then via the generative modeling techniques, we would be able to fill the archive successfully without being limited by the run times of the evolutionary algorithm and its sequential nature. Another interesting direction would be an epistemic uncertainty-guided evolution frame. The intuition here lies in the fact that epistemic uncertainty decreases when the number of data we train our model increases [hullermeier2021aleatoric], hence, we can use the epistemic uncertainty of the model as a way to guide the QD evolution into more uncertain behaviors, with the goal of reducing the epistemic uncertainty of the VQ-VAE model, and consequently creating the most optimal behavioral space possible. For this purpose, we can borrow ideas from surrogate-assisted QD algorithms [kent2024bayesian]. Lastly, we can combine the unsupervised QD with elements of information theory, specifically the normalized entropy [wilcox1967indices], to devise a more concrete metric to assess the quality of the behavior space.